# YITP-26-12, RESCEU-4/26, IPMU26-0003

### v1

@@ II.5 Simple example
The result of the procedure described in § II.4 is summarized on Fig. 2 that shows $H_{\rm E}^{2}$, $a$, $\dot{\psi}$ and $\dot{H}_{\rm E}$ as functions of $\psi$ and then $\psi$ and $H_{E}$ as functions of $z$, where we have set $N_{\rm E}=1$. The profile of the clock field allows, depending on the mass scale $M$, for the apparition of two Lorentzian universes respectively for $z>z_{c}$ and $z<-z_{c}$ with $z_{c}$ the positive root of Eq. (19). Then at large $z$, $\psi$ tends to a constant so that $t\sim z$ and since $H_{E}$ also tends to a constant, $H$ tends to a constant at large $t$. Hence, this conforts us to conclude that two mirror emergent universes with a proper primordial almost de Sitter phase and “emergent time flowing in different directions” can be obtained with the simple choice of the free functions.
@@ III.1 Formalism
$\mathcal{K}(X_{\rm E})=6H_{E}^{2}\mathcal{G}(X_{\rm E})+2J_{0}\frac{\sqrt{X_{\rm E}}}{a^{3}}\,$, (30)
@@
Let us also stress that the case $J_{0}=0$, in which the formal integrals simplify considerably, must be handled with care. In this case one reconstructs $\mathcal{G}(X_{\rm E})=C_{0}/\tilde{H}_{\rm E}(X_{\rm E})$ and $\mathcal{K}(X_{\rm E})=6C_{0}\tilde{H}_{\rm E}(X_{\rm E})$, where we have renamed the input $H_{\rm E}(X_{\rm E})$ as $\tilde{H}_{\rm E}(X_{\rm E})$. However, once the Lagrangian is reconstructed in this way, the set of equations (13) for $\{h(z)$, $\phi(z)\}$ generically becomes ill-behaved: (i) if $\tilde{H}_{\rm E}$ is not proportional to $\sqrt{X_{\rm E}}$ and if the algebric equation $2\tilde{H}^{\prime}_{\rm E}(X_{\rm E})X_{\rm E}=\tilde{H}_{\rm E}(X_{\rm E})$ does not admit a non-negative real solution then there is no solution with $J_{0}\neq 0$; (ii) if $\tilde{H}_{\rm E}$ is not proportional to $\sqrt{X_{\rm E}}$ and if the algebraic equation admits a non-negative real solution $X_{\rm E}=X_{{\rm E}*}=const.$ then the only solution with $J_{0}\neq 0$ is those with $\psi(z)=\pm\sqrt{X_{{\rm E}*}}$ and thus generically exhibits discontinuity at $J_{0}=0$ where $\psi(z)$ is undetermined by the equations of motion; (ii) if $\tilde{H}_{\rm E}$ is proportional to $\sqrt{X_{\rm E}}$ then $\psi(z)$ is undetermined by the equations of motion.
@@ III.2 Worked out example and implications
which has the property that $h\sim kz$ at large $|z|$ so that $H_{\rm E}(z)\rightarrow\pm k$. Then, two kinds of configurations can be constructed depending on the clock field profile, either mirror universes, as already obtained, or pocket universes embedded in a Euclidean sea, a simple choice being
@@
with $\Delta\psi<\psi_{0}$ and $\psi_{0}>0$ to ensure that $\psi$ does not cross 0, as imposed by the analysis of § II.2. Figure 3 illustrates these two classes of models and Fig. 4 depicts the reconstructed free functions $\{{\cal K}(X_{\rm E}),G_{4}(X_{\rm E})\}$ that depend on the two integration constants $(J_{0},C_{0})$ assuming, as discussed above, that $C_{1}=0$. To finish, we give the emergent cosmologies, that is $R(t)$ and $H(t)$ in Fig. 5.
@@
This demonstrates that there exist solutions of the two families. Concerning the mirror universes, the expansion starts from a finite value of the scale factor at $t=0$ on $\Sigma_{0}$ and expands indefinitely while converging to a de Sitter behavior. Interestingly, due to the variation of $\psi$ close to $\Sigma_{0}$, the universe has a primordial phase prior to its de Sitter phase. This can eventually have cosmological signatures the study of which is beyond the scope of this letter. Concerning the pocket universes, it has a finite lifetime since there is a boundary $\Sigma_{0}$ in the future. Since $\psi$ has no time to settle to constant value, the scale factor never behaves as de Sitter but enjoys a mild bounce.
Figure 4: Reconstruction of the free functions ${\cal K}(X_{\rm E})$ [Top] and $G_{4}(X_{\rm E})$ [Bottom] for a mirror universes [Solid lines] and pocket universe [Dashed lines] assuming $J_{0}=1$ and $C_{0}=0$. The mirror universe curve is not defined below the value of $X_{E}$ corresponding to the minimum value of $\psi$ occurring at $z=0$ in Fig. 3, and from the form of Eq. (35), one can see that the values of $\psi$ and consequently $X_{E}$ are bounded.
Figure 5: The emergent cosmologies for mirror universes [Solid lines] and pocket universe [Dashed lines]. The scale factor $R$ and the Hubble rate $H$ as a function of the cosmic time $t$ – assuming $t=0$ on $\Sigma_{0}$ – are respectively plotted on the top and bottom frames for the same parameters as in Fig. 4. The red dot-dashed curve represents an exact de Sitter spacetime and the vertical dashed lines mark the limit of the Lorentzian regions ${\cal M}_{0}$. Contrary to the mirror universes, a pocket universe has a boundary in the future and hence a finite maximum age.
@@ IV Discussion
This letter gives a proof of concept that within the framework for the emergence of the Lorentz signature originally proposed by Refs. Mukohyama:2013ew ; Kehayias:2014uta , one can construct pocket universes with emergent dynamical expansion of Friedmann-Lemaître kind. In this approach, the standard big bang singularity is replaced by a 3 dimensional hypersurface $\Sigma_{0}$ on which the signature of the metric flips. Hence, dynamics freezes outside the Lorentzian pockets that eventually emerge. It should be emphasized that the Euclidean metric $g^{\rm E}_{\mu\nu}$ and the scalar clock field $\phi$ remain regular at $\Sigma_{0}$, so it is not a singularity in these variables.
Our analysis relies on the action (7) initially introduced in Ref. Mukohyama:2013ew . Our construction also assumed a planar symmetry in order to recover a spatially homogeneous and isotropic spacetime in the Lorentzian regions, as depicted on Fig. 1. But this can only be thought as a local approximation since this symmetry can a priori not be extended over the whole 4-dimensional space. Hence, our universe may also have a spatial boundary on scale larger than our observable universe.
Figure 6: Once boundary conditions are set, the profile of the clock field cannot cross the line $\psi=0$ as demonstrated in § II.2. $\psi_{c}=M^{2}$ defines the change-of-signature hypersurface $\Sigma_{0}$ and delimitates the Lorentzian and Euclidean regions; see Eq. (19). $\psi_{0}$ is defined by the condition $j(\psi_{0})=0$; see Eq. (9). From a generic point of view, we may expect to have a non monotonic profile so that the 4 dimensional- Riemannian space may contain many Lorentzian “pockets” that may or not be bounded in the $z$, i.e. emergent time, direction.
@@
First, we have derived in § II.2 a necessary condition for a Lorentzian pocket to emerge, in the form of a no-go theorem. We have then studied in § II.4 the cases that evade the no-go and showed an explicit example in which a Euclidean island is embedded between universes with the Lorentzian signature. In this case the emergent geometry in each of the two asymptotic regions enjoys an almost de Sitter expansion while the initial cosmological singularity of each universe is replaced by the hypersurface $\Sigma_{0}$ on which the effective metric becomes degenerate and beyond which it has the signature $(+,+,+,+)$. Then § III has demonstrated that the reconstruction procedure can be explicitly solved, i.e., one can determine the two free functions $G_{4}$ and $\cal K$} from the choice of the clock field and spatial geometry profiles, that is from $\psi(z)$ and $a(z)$. This approach allowed us to show that our model contains a rich phenomenology containing at least to classes of emergent spacetime: mirror universes already discussed in § II.4 and pocket universes; see Fig. 4. In the latter case, the Friedmann-Lemaître spacetime exhibits a bounce but the dynamics is bounded by $\Sigma_{0}$ both in the future and the past; see Fig. 5.
While beyond the scope of this first analysis, it is clear that in principle one could mimic the usual cosmological big bang expansion history $a_{\rm BB}(t)$ after a choice of clock field profile $\psi(z)$. Indeed, imagine that $\psi$ becomes linear outside of a neighborhood of $\Sigma_{0}$$\psi(|z|>z_{0})\sim M|z|$ – then Eq. (23) is integrable as
@@
from which it follows that $h(z)$ is given by $\ln a_{\rm BB}[t(z)]$ and the reconstruction procedure ensures that in this ad hoc setting, a model reproducing the whole expansion history of our observable universe can be constructed. Indeed, such an argument of existence is far from being sufficient since we shall consider free functions that are fixed and study e.g., whether this requires a fine tuning and the behaviour of solutions with other symmetries, etc.
Even if this new cosmological construction may seem very speculative, it offers a new vision of the primordial universe and of the universe on scales larger than those of our observable universe, which gives a concrete reconstruction of the early insight by Sakharov Sakharov:1984csx . Besides, it has specific properties and signatures. First, the big bang is no more a singularity in the fundamental (Euclidean) variables and $R(t=0)$ remains finite as seen on Fig. 3. Second, the expansion in the primordial phase may not be directly de Sitter so that inflation would be preceded by an expansion phase that could let a specific signature on the primordial fluctuations. Indeed, this remains to be investigated with care. Third, the physical extension of the 3-dimensional spatial (i.e. constant $t$ or $z$) hypersurfaces may be bounded so that one may have a finite universe without invoking a spatial topology. Indeed it has to be larger than the observable universe but may still leave detectable signatures, e.g. on CMB temperature and polarisation distributions Fabre:2013wia . To finish, there shall be gravity beyond General Relativity and the amplitude of the deviations from GR can be predicted once a full model is constructed and its parameters determined by the cosmology, hence making the construction, at least in principle, falsifiable. The propagation of test fields in spacetime with Euclidean pockets has already been studied in analogue systems Weinfurtner:2007br ; Donley:2001wvu ; Weinfurtner:2007dq , which can be a setting for exploring primordial fluctuations.
It is also worth pointing out that the change in the metric signature of the effective metric $g_{\mu\nu}$ fundamentally modifies the nature of the matter field equations since they shift from elliptic to hyperbolic – as was actually the early motivation of Refs. Mukohyama:2013ew ; Kehayias:2014uta ; Mukoyama:2013gqu ; Mukohyama:2013gra ; Uzan-CUP . While a Lorentzian effective metric is required at long distance scales, on short distance scales in the vicinity of $\Sigma_{0}$, one will need to introduce higher derivative terms in the matter action coupled to $g^{\rm E}_{\mu\nu}$ in order to avoid the potential pathologies that arise from coupling to a signature changing metric. In a similar manner, a more complete treatment of the gravitational sector will require a higher derivative model as well, following that introduced in Ref. Mukohyama:2013gra and Eq. (2.5) therein, i.e.,
@@
To conclude, we can grasp the big picture on the universe at large that emerges when the description developed in this Letter. It is illustrated in Figs. 6 and 7. Since the large scale distribution of the clock field shall depend on boundary conditions and eventually on the geometry of the 4-dimensional Euclidean space, one expects it to be non-trivial. In full generalities, many Lorentzian pockets can emerge with their own geometry and with a large span of sizes, i.e. of expansion time as seen form their inside. It also follows that the description of the primordial universe may not require to go through a quantum gravity phase even if it requires gravity beyond General Relativity. This proposal opens a new, rich and exciting phenomenology that needs to be investigated further and that can eventually change our conception of the cosmos on the largest scales.
@@ Acknowledgements.
We are deeply grateful to Sante Carloni for helpful discussions during the early stages of this work. J. C. F. is grateful for the hospitality of the Research Center for the Early Universe (RESCEU) and University of Tokyo, as well as the Yukawa Institute for Theoretical Physics (YITP), Kyoto University, during which part of this work was performed. J. C. F. is supported by the European Union and Czech Ministry of Education, Youth and Sports through the FORTE project No. CZ.02.01.01/00/22_008/0004632. The work of S. M. was supported in part by JSPS (Japan Society for the Promotion of Science) KAKENHI Grant No. JP24K07017 and World Premier International Research Center Initiative (WPI), MEXT, Japan. J. P. U. is thankful to YITP for hospitality during his visit in fall 2025.

### v2

@@ II.5 Simple example
The result of the procedure described in Sec. II.4 is summarized in Fig. 2 that shows $H_{\rm E}^{2}$, $a$, $\dot{\psi}$ and $\dot{H}_{\rm E}$ as functions of $\psi$ and then $\psi$ and $H_{E}$ as functions of $z$, where we have set $N_{\rm E}=1$. The profile of the clock field allows, depending on the mass scale $M$, for the apparition of two Lorentzian universes respectively for $z>z_{c}$ and $z<-z_{c}$ with $z_{c}$ the positive root of Eq. (19). Then at large $z$, $\psi$ tends to a constant so that $t\sim z$ and since $H_{E}$ also tends to a constant, $H$ tends to a constant at large $t$. Hence, this allows us to conclude that two mirror emergent universes with a proper primordial almost de Sitter phase and “emergent time flowing in different directions” can be obtained with the simple choice of the free functions.
@@ III.1 Formalism
$\mathcal{K}(X_{\rm E})=6H_{E}^{2}\mathcal{G}(X_{\rm E})+2J_{0}\frac{\sqrt{X_{\rm E}}}{a^{3}}\,$ (30)
@@
Let us also stress that the case $J_{0}=0$, in which the formal integrals simplify considerably, must be handled with care. In this case one reconstructs $\mathcal{G}(X_{\rm E})=C_{0}/\tilde{H}_{\rm E}(X_{\rm E})$ and $\mathcal{K}(X_{\rm E})=6C_{0}\tilde{H}_{\rm E}(X_{\rm E})$, where we have renamed the input $H_{\rm E}(X_{\rm E})$ as $\tilde{H}_{\rm E}(X_{\rm E})$. However, once the Lagrangian is reconstructed in this way, the set of equations (13) for $\{h(z)$, $\phi(z)\}$ generically becomes ill behaved: (i) if $\tilde{H}_{\rm E}$ is not proportional to $\sqrt{X_{\rm E}}$ and if the algebraic equation $2\tilde{H}^{\prime}_{\rm E}(X_{\rm E})X_{\rm E}=\tilde{H}_{\rm E}(X_{\rm E})$ does not admit a non-negative real solution, then there is no solution with $J_{0}\neq 0$; (ii) if $\tilde{H}_{\rm E}$ is not proportional to $\sqrt{X_{\rm E}}$ and if the algebraic equation admits a non-negative real solution $X_{\rm E}=X_{{\rm E}*}=const.$ then the only solution with $J_{0}\neq 0$ is those with $\psi(z)=\pm\sqrt{X_{{\rm E}*}}$ and thus generically exhibits discontinuity at $J_{0}=0$ where $\psi(z)$ is undetermined by the equations of motion; (iii) if $\tilde{H}_{\rm E}$ is proportional to $\sqrt{X_{\rm E}}$ then $\psi(z)$ is undetermined by the equations of motion.
@@ III.2 Worked out example and implications
which has the property that $h\sim kz$ at large $|z|$ so that $H_{\rm E}(z)\rightarrow\pm k$. Then, two kinds of configurations can be constructed depending on the clock-field profile, either mirror universes, as already obtained, or pocket universes embedded in a Euclidean sea, a simple choice being
@@
with $\Delta\psi<\psi_{0}$ and $\psi_{0}>0$ to ensure that $\psi$ does not cross 0, as imposed by the analysis of Sec. II.2. Figure 3 illustrates these two classes of models and Fig. 4 depicts the reconstructed free functions $\{{\cal K}(X_{\rm E}),G_{4}(X_{\rm E})\}$ that depend on the two integration constants $(J_{0},C_{0})$ assuming, as discussed above, that $C_{1}=0$. To finish, we give the emergent cosmologies, that is $R(t)$ and $H(t)$ in Fig. 5.
@@
This demonstrates that there exist solutions of the two families. Concerning the mirror universes, the expansion starts from a finite value of the scale factor at $t=0$ on $\Sigma_{0}$ and expands indefinitely while converging to a de Sitter behavior. Interestingly, due to the variation of $\psi$ close to $\Sigma_{0}$, the universe has a primordial phase prior to its de Sitter phase. This can eventually have cosmological signatures the study of which is beyond the scope of this paper. Concerning the pocket universes, it has a finite lifetime since there is a boundary $\Sigma_{0}$ in the future. Since $\psi$ has no time to settle to constant value, the scale factor never behaves as de Sitter but enjoys a mild bounce.
Figure 4: Reconstruction of the free functions ${\cal K}(X_{\rm E})$ [top] and $G_{4}(X_{\rm E})$ [bottom] for a mirror universe [solid lines] and pocket universe [dashed lines] assuming $J_{0}=1$ and $C_{0}=0$. The mirror universe curve is not defined below the value of $X_{E}$ corresponding to the minimum value of $\psi$ occurring at $z=0$ in Fig. 3, and from the form of Eq. (35), one can see that the values of $\psi$ and consequently $X_{E}$ are bounded.
Figure 5: The emergent cosmologies for mirror universes [solid lines] and pocket universe [dashed lines]. The scale factor $R$ and the Hubble rate $H$ as a function of the cosmic time tt—assuming $t=0$ on $\Sigma_{0}$—are respectively plotted on the top and bottom frames for the same parameters as in Fig. 4. The red dot-dashed curve represents an exact de Sitter spacetime and the vertical dashed lines mark the limit of the Lorentzian regions ${\cal M}_{0}$. Contrary to the mirror universes, a pocket universe has a boundary in the future and hence a finite maximum age.
@@ IV Discussion
This paper gives a proof of concept that within the framework for the emergence of the Lorentz signature originally proposed by Refs. Mukohyama:2013ew ; Kehayias:2014uta , one can construct pocket universes with emergent dynamical expansion of Friedmann-Lemaître kind. In this approach, the standard big bang singularity is replaced by a three dimensional hypersurface $\Sigma_{0}$ on which the signature of the metric flips. Hence, dynamics freezes outside the Lorentzian pockets that eventually emerge. It should be emphasized that the Euclidean metric $g^{\rm E}_{\mu\nu}$ and the scalar clock field $\phi$ remain regular at $\Sigma_{0}$, so it is not a singularity in these variables.
Our analysis relies on the action (7) initially introduced in Ref. Mukohyama:2013ew . Our construction also assumed a planar symmetry in order to recover a spatially homogeneous and isotropic spacetime in the Lorentzian regions, as depicted in Fig. 1. But this can only be thought as a local approximation since this symmetry can a priori not be extended over the whole four-dimensional space. Hence, our Universe may also have a spatial boundary on scale larger than our observable Universe.
Figure 6: Once boundary conditions are set, the profile of the clock field cannot cross the line $\psi=0$ as demonstrated in Sec. II.2. $\psi_{c}=M^{2}$ defines the change-of-signature hypersurface $\Sigma_{0}$ and delimitates the Lorentzian and Euclidean regions; see Eq. (19). $\psi_{0}$ is defined by the condition $j(\psi_{0})=0$; see Eq. (9). From a generic point of view, we may expect to have a nonmonotonic profile so that the four dimensional- Riemannian space may contain many Lorentzian “pockets” that may or not be bounded in the $z$, i.e. emergent time, direction.
@@
First, we have derived in Sec. II.2 a necessary condition for a Lorentzian pocket to emerge, in the form of a no-go theorem. We have then studied in Sec. II.4 the cases that evade the no-go and showed an explicit example in which a Euclidean island is embedded between universes with the Lorentzian signature. In this case the emergent geometry in each of the two asymptotic regions enjoys an almost de Sitter expansion while the initial cosmological singularity of each universe is replaced by the hypersurface $\Sigma_{0}$ on which the effective metric becomes degenerate and beyond which it has the signature $(+,+,+,+)$. Then Sec. III has demonstrated that the reconstruction procedure can be explicitly solved, i.e., one can determine the two free functions $G_{4}$ and $\cal K$} from the choice of the clock field and spatial geometry profiles, that is from $\psi(z)$ and $a(z)$. This approach allowed us to show that our model contains a rich phenomenology containing at least to classes of emergent spacetime: mirror universes already discussed in Sec. II.4 and pocket universes; see Fig. 4. In the latter case, the Friedmann-Lemaître spacetime exhibits a bounce but the dynamics is bounded by $\Sigma_{0}$ both in the future and the past; see Fig. 5.
While beyond the scope of this first analysis, it is clear that in principle one could mimic the usual cosmological big bang expansion history $a_{\rm BB}(t)$ after a choice of clock-field profile $\psi(z)$. Indeed, imagine that $\psi$ becomes linear outside of a neighborhood of $\Sigma_{0}$$\psi(|z|>z_{0})\sim M|z|$ – then Eq. (23) is integrable as
@@
from which it follows that $h(z)$ is given by $\ln a_{\rm BB}[t(z)]$ and the reconstruction procedure ensures that in this ad hoc setting, a model reproducing the whole expansion history of our observable Universe can be constructed. Indeed, such an argument of existence is far from being sufficient since we shall consider free functions that are fixed and study e.g., whether this requires a fine-tuning and the behavior of solutions with other symmetries, etc.
Even if this new cosmological construction may seem very speculative, it offers a new vision of the primordial Universe and of the Universe on scales larger than those of our observable Universe, which gives a concrete reconstruction of the early insight by Sakharov Sakharov:1984csx . Besides, it has specific properties and signatures. First, the big bang is no more a singularity in the fundamental (Euclidean) variables and $R(t=0)$ remains finite as seen in Fig. 3. Second, the expansion in the primordial phase may not be directly de Sitter so that inflation would be preceded by an expansion phase that could let a specific signature on the primordial fluctuations. Indeed, this remains to be investigated with care. Third, the physical extension of the three-dimensional spatial (i.e., constant $t$ or $z$) hypersurfaces may be bounded so that one may have a finite universe without invoking a spatial topology. Indeed it has to be larger than the observable Universe but may still leave detectable signatures, e.g. on CMB temperature and polarization distributions Fabre:2013wia . To finish, there shall be gravity beyond general relativity and the amplitude of the deviations from general relativity can be predicted once a full model is constructed and its parameters determined by the cosmology, hence making the construction, at least in principle, falsifiable. The propagation of test fields in spacetime with Euclidean pockets has already been studied in analogue systems Weinfurtner:2007br ; Donley:2001wvu ; Weinfurtner:2007dq ; Nissinen:2017 , which can be a setting for exploring primordial fluctuations.
It is also worth pointing out that the change in the metric signature of the effective metric $g_{\mu\nu}$ fundamentally modifies the nature of the matter field equations since they shift from elliptic to hyperbolic—as was actually the early motivation of Refs. Mukohyama:2013ew ; Kehayias:2014uta ; Mukoyama:2013gqu ; Mukohyama:2013gra ; Uzan-CUP . While a Lorentzian effective metric is required at long distance scales, on short distance scales in the vicinity of $\Sigma_{0}$, one will need to introduce higher derivative terms in the matter action coupled to $g^{\rm E}_{\mu\nu}$ in order to avoid the potential pathologies that arise from coupling to a signature changing metric. In a similar manner, a more complete treatment of the gravitational sector will require a higher derivative model as well, following that introduced in Ref. Mukohyama:2013gra and Eq. (2.5) therein, i.e.,
@@
To conclude, we can grasp the big picture on the Universe at large that emerges when the description developed in this paper. It is illustrated in Figs. 6 and 7. Since the large-scale distribution of the clock field shall depend on boundary conditions and eventually on the geometry of the four-dimensional Euclidean space, one expects it to be nontrivial. In full generalities, many Lorentzian pockets can emerge with their own geometry and with a large span of sizes, i.e. of expansion time as seen from their inside. It also follows that the description of the primordial Universe may not require it to go through a quantum gravity phase even if it requires gravity beyond general relativity. This proposal opens a new, rich and exciting phenomenology that needs to be investigated further and that can eventually change our conception of the cosmos on the largest scales.
@@ Acknowledgements.
We are deeply grateful to Sante Carloni for helpful discussions during the early stages of this work. J. C. F. is grateful for the hospitality of the Research Center for the Early Universe (RESCEU) and University of Tokyo, as well as the Yukawa Institute for Theoretical Physics (YITP), Kyoto University, during which part of this work was performed. J. C. F. is supported by the European Union and Czech Ministry of Education, Youth and Sports through the FORTE Project No. CZ.02.01.01/00/22_008/0004632. The work of S. M. was supported in part by JSPS (Japan Society for the Promotion of Science) KAKENHI Grant No. JP24K07017 and World Premier International Research Center Initiative (WPI), MEXT, Japan. J.-P. U. is thankful to YITP for hospitality during his visit. The Mathematica file used to perform the numerical integration is available in the GitHub repository Feng:2026mat .